# Discrete homotopy and homology theories for finite posets

### v1

@@ 1. Introduction
The classical homotopy and homology theories for finite posets are based on a well-known correspondence given by McCord, which assigns to each finite poset $M$ the order complex $\mathcal{K}(M)$ together with a weak homotopy homotopy equivalence $|\mathcal{K}(M)|\rightarrow M$ [1, Theorem 1.4.6]. This correspondence permits the study of homotopy and homology groups of a finite poset by means of its associated order complex.
@@
The motivation for the second result comes from a discrete cubical homology theory for graphs introduced in [3, 4]. We define a similar discrete cubical homology theory for finite posets. Our second result is that we exhibit for each $n\geq 1$ a natural homomorphism $\psi:H_{n}^{\rm Cube}(M)\rightarrow H_{n}^{\rm Simpl}(\mathcal{K}(M))$, where $H_{n}^{\rm Cube}(M)$ is the $n$th discrete cubical homology group, and a class of finite posets $M$ for which $\psi$ is surjective (Theorem 4.7). We provide an example for which the homology groups of the two theories are not isomorphic at least in dimension $1$ (Example 4.9).
Theorem 4.7 has an interesting application as follows. Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible. Corollary 4.17 tells when the manifold $M$ with two specific balls removed is not contractible.
@@
In this paper, all maps between posets that we deal with are assumed to be order-preserving.
@@ 2. Preliminaries
This section is devoted to introducing some definitions and properties of finite posets and simplicial complexes that are needed in the sequel. We refer the reader to [1, 7, 11, 12] for further details or proofs.
@@ 2.1. Finite posets
A topological space is said to be Alexandroff if every intersection of open sets is still an open set. If $X$ is an Alexandroff space, for each $x\in X,$ we denote by $U_{x}$ the intersection of all the open sets of $X$ containing $x$. This allows us to define a reflexive and transitive relation on $X$ by declaring $x\leq y$ if $U_{x}\subseteq U_{y}.$ The relation is antisymmetric if and only if $X$ is $T_{0}$. On the other hand, if $X$ is a set endowed with a reflexive, transitive relation, then the sets $U_{x}=\{y\in X|y\leq x\}$, as a basis, determines an Alexandroff topology. The Alenxandroff topology is $T_{0}$ if and only if the relation is antisymmetric, that is, a partial order. A set with a partial order is called a poset. It has been shown that Alexandroff $T_{0}$ spaces and posets are in bijective correspondence. A finite $T_{0}$ space is clearly an Alexandroff $T_{0}$ space, thus a finite $T_{0}$ space is equivalent to a finite poset.
@@
Given a poset $X$ and an element $x\in X$, we denote $U_{x}=\{y\in X~{}|~{}y\leq x\},~{}\widehat{U}_{x}=\{y\in X~{}|~{}y<x\},$ $F_{x}=\{y\in X~{}|~{}x\leq y\},$ and $\widehat{F}_{x}=\{y\in X~{}|~{}x<y\}.$ Let $C_{x}=U_{x}\cup F_{x}$ and $\widehat{C}_{x}=\widehat{U}_{x}\cup\widehat{F}_{x}$. A totally ordered subset $C$ of $X$ is called a chain. An $n$-chain of $X$ is a chain of cardinality of $n+1$.
@@
A finite poset $X$ is homogeneous of dimension $n$ if all maximal chains in $X$ have cardinality of $n+1$. In this case, the degree of $x$, denoted ${\rm deg}(x)$, is the dimension of $U_{x}$.
If $X$ is a finite poset, we define its order complex $\mathcal{K}(X)$ as the ordered simplicial complex whose simplices are the nonempty chains of $X$. Conversely, given a finite simplicial complex $K$, we define the face poset $\chi(K)$ of $K$ as the poset of simplices of $K$ ordered by inclusion.
@@
A result of McCord [1, Theorem 1.4.6] asserts that $\mu_{X}$ is a weak homotopy equivalence. In fact, the same argument in [1, Theorem 1.4.6] shows $\widehat{\mu_{X}}:|\mathcal{K}(X)|\rightarrow X$ defined by
(2.1) $\widehat{\mu_{X}}(u)=\text{max}(\text{support}(u))$
is also a weak homotopy equivalence.
@@ 2.2. Simplicial complexes
Let $K$ be a simplicial complex. A simplex $A\subseteq K$ is called a principle simplex if $A$ is not a proper face of any simplex of $K$. The face $B$ of $A$ is called a free face of $A$ in $K$ if $B$ is the proper face of no other simplex of $K$.
$K$ is a principle $k$-simplicial complex if every principle simplex of $K$ is $k$-dimensional.
@@
Suppose $X\subseteq Y$ are polyhedra and that $Y=X\cup B^{n}$ and $X\cap B^{n}=$ a face $B^{n-1}$, where $B^{i}$ denotes a piecewise-linear (PL) $i$-ball. Then we say that there is an elementary collapse of $Y$ onto $X$, and write $Y\searrow X$. The collapse is across $B^{n}$ onto $B^{n-1}$ from the complementary face ${\rm cl}(\partial B^{n}-B^{n-1})$.
We say $Y$ collapses to $X$, writen $Y\searrow X$, if there is a finite sequence of elementary collapses $Y=Y_{0}\searrow Y_{1}\searrow\cdots\searrow Y_{n}=X$. If $X$ is a point, we call $Y$ collapsible and write $Y\searrow 0$.
@@
3. Discrete homotopy theory for fintie posets
@@
Let $\mathbb{Z}$ denote the digraph whose vertices are integers and whose edges are the ordered pairs $(2i,2i+1)$ and $(2j,2j-1)$. Given an integer $p\geq 0$, write $I_{p}$ for the subgraph of $\mathbb{Z}$ induced by $\{0,\cdots,p\}$.
@@
Let $\mathbb{Z}^{n}$ be the $n$-fold Cartesian product of $\mathbb{Z}$ whose vertices are $n$-tuples of integers. For any $r\geq 0$, $\mathbb{Z}^{n}_{\geq r}$ denotes the subgraph consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}$ with $|x_{i}|=r$ for some $i$, and $\mathbb{Z}^{n}_{\leq r}$ denotes the subgraph consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}$ with $|x_{i}|\leq r$ for every $i$. Denote by $\partial\mathbb{Z}^{n}_{\leq r}$ the subgraph of $\mathbb{Z}^{n}_{\leq r}$ consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}_{\leq r}$ with $|x_{i}|=r$ for some $i$. Similarly, Denote by $\partial I^{n}_{\leq r}$ the subgraph of $I^{n}_{\leq r}$ consisting of all vertices $x=(x_{1},\cdots,x_{n})\in I^{n}_{\leq r}$ with $x_{i}=r$ or 00 for some $i$.
@@
if there exists an $r\geq 0$ such that $f$ is a map $f:(\mathbb{Z}^{n},\mathbb{Z}_{\geq r}^{n})\rightarrow(X,x_{0})$. The minimum of such $r$ is called the radius of $f$
@@
For a finite poset $X$ with basepoint $x_{0}\in X$, define $\pi_{n}^{D}(X,x_{0})$ to be the set of basepoint-preserving-homotopy classes of maps $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$. We call $\pi_{n}^{D}(X,x_{0})$ the $n$th discrete homotopy group of $X$.
@@
Let $L^{n}$ be the $n$-dimensional cube, the product of $n$ copies of the interval $[-1,1]$. Given a partition
(3.1) $-1=t_{-\widehat{m_{i}}}<t_{-\widehat{m_{i}}+1}<\cdots<t_{-1}<0=t_{0}<t_{1}\cdots<t_{m_{i}-1}<t_{m_{i}}=1,$
of the $i$-th copy of $L^{n}$ for each $i$, $\widehat{m_{i}}$ being a positive integer, $L^{n}$ can be identified with the geometric realization of a subgraph $N$ of $\mathbb{Z}^{n}$,
$N=\prod_{i}^{n}\mathbb{Z}_{[-\widehat{m_{i}},m_{i}]}.$
Write $P$ for this partition. The subgraph $N$ is denoted $\gamma_{P}$. For each vertex $(x_{1},x_{2},\cdots,x_{n})$ of $\gamma_{P}$, taking it to the point $(t_{x_{1}},t_{x_{2}},\cdots,t_{x_{n}})$ of $L^{n}$ induces a canonical homeomorphism
$\lambda_{P}:|\mathcal{K}(\gamma_{P})|$ $\longrightarrow L^{n}.$
When the partition $P$ takes the form
$-1=\frac{-r}{r}<\frac{-r+1}{r}<\cdots<\frac{-1}{r}<0=t_{0}<\frac{1}{r}\cdots<\frac{r-1}{r}<\frac{r}{r}=1$
for each $i$, we will write $P$, $\gamma_{P}$ and $\lambda_{P}$ as $P_{r}$, $\gamma_{r}$ and $\lambda_{r}$, respectively.
Let $I$ be the unit interval $[0,1]$. The geometric realization of the subgraph, $I_{q}$, of $\mathbb{Z}$ can be viewed as $I$ by the canonical homeomorphism
$\beta_{q}:|\mathcal{K}(I_{q})|\rightarrow I$
with $j$ mapping to $\frac{j}{q}$ for each vertex $j$.
@@
It is straightforward to check that $\theta$ is well-defined and is a homomorphism.
@@
Since $X$ is connected, we may assume that $x_{0}$ is a minimum vertex of $X$. Suppose $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ is a map of radius $r$ such that $\theta([f])=0$. Let $G:(L^{n},\partial L^{n})\times I\rightarrow(|\mathcal{K}(X)|,x_{0})$ be a homotopy from $|\mathcal{K}(f)|\circ\lambda_{r}^{-1}$ to the constant map. Note that $G$ can be written as
$G:\left(\lambda_{r}(|\mathcal{K}(\gamma_{r})|),\partial\lambda_{r}(|\mathcal{K}(\gamma_{r})|)\right)\times|\mathcal{K}(I_{q})|\rightarrow(|\mathcal{K}(X)|,x_{0}).$
Cover $\lambda_{r}(|\mathcal{K}(\gamma_{r})|)\times|\mathcal{K}(I_{q})|$ by the open sets
$(\widehat{\mu_{X}}\circ G)^{-1}(U^{X}_{u}),$
as $u$ ranges over all vertices of $X$. Denote this open covers by $\mathcal{A}$.
Let $\delta_{\mathcal{A}}$ be the Lebesgue number for the open cover $\mathcal{A}$. We take $r_{1}$ to be a multiple of $r$ and $q_{1}$ to be a even integer larger than $q$ so that each subpolyhedron of $\lambda_{r_{1}}(|\mathcal{K}(\gamma_{r_{1}})|)$, which is the image of a simplex, has diameter less than $\frac{\delta}{2}$, where $\delta<\delta_{\mathcal{A}}$. Let us construct a homotopy
$H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{q_{1}}\rightarrow(X,x_{0})$
as follows. If $(w,t)\in\mathbb{Z}_{\leq r_{1}}^{n}\times I_{q_{1}}-\mathbb{Z}_{\leq r_{1}}^{n}\times 0-\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$, then we have
(3.2) $\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t})|)\subseteq(\widehat{\mu_{X}}\circ G)^{-1}(U^{X}_{u}),$
for some vertex $u$ of $X$, $\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times|\mathcal{K}(U^{I_{q_{1}}}_{t})|$ being a subset of the closed star of $(\lambda_{r_{1}}(w),t)$. In this case, set
$H(w,t)=u,$
@@
(3.3) $H(w,t)=\begin{cases}x_{0}~{}&\text{if}~{}w\in\partial\mathbb{Z}_{\leq r_{1}}^{n}~{}\text{or}~{}t=q_{1},\\ \widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)~{}&\text{if}~{}(w,t)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 0.\end{cases}$
Let us check that $H$ is an order-preserving. We only have to consider three cases:
@@
    $(w,t),(w^{\prime},t^{\prime})\in\mathbb{Z}_{\leq r_{1}}^{n}\times I_{q_{1}}-\mathbb{Z}_{\leq r_{1}}^{n}\times 0-\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$ and $(w,t)\leq(w^{\prime},t^{\prime})$.

### v2

@@ 1. Introduction
The classical homotopy and homology theories for finite posets are based on a well-known correspondence given by McCord, which assigns to each finite poset $X$ the order complex $\mathcal{K}(X)$ together with a weak homotopy equivalence $|\mathcal{K}(X)|\rightarrow X$ [1, Theorem 1.4.6]. This correspondence permits the study of homotopy and homology groups of a finite poset by means of its associated order complex.
The classical homotopy-theoretic invariants for a finite poset involve only its topological structure, that is, the associated order complex. However, a salient feature of a finite poset is that it is essentially a combinatorial object, its graphical representation (Hasse diagram) being a digraph. The combinatorial properties of posets have been applied to other areas, such as topological data analysis. For example, in [12] Jardine utilizes poset theoretical ideas to study Vietoris-Rips complexes associated with metric spaces. Moreover, in light of the observation that every finite digraph can be viewed as the graphical representation (Hasse diagram) of a finite poset, the study of the combinatorics of posets can facilitate the understanding of the combinatorics of digraphs. The homotopy-theoretic invariants [5, 6], based on combinatorics of the latter, have found numerous applications, including in subspace arrangements [7], pattern recognition [8] and topological data analysis [13]. Hence, the motivation for the introduction of a new discrete homotopy theory for finite posets in this paper comes from a desire to find invariants that are sensitive to the combinatorics encoded in the posets.
@@
The motivation for the second result comes from a discrete cubical homology theory for graphs introduced in [4, 3]. We define a similar discrete cubical homology theory for finite posets. Our second result is that we exhibit for each $n\geq 1$ a natural homomorphism $\psi:H_{n}^{\rm Cube}(X)\rightarrow H_{n}^{\rm Simpl}(\mathcal{K}(X))$, where $H_{n}^{\rm Cube}(X)$ is the $n$-th discrete cubical homology group, and a class of finite posets $X$ for which $\psi$ is surjective (Theorem 4.7). We provide an example for which the homology groups of the two theories are not isomorphic at least in dimension $1$ (Example 4.10). Theorem 4.7 has an interesting application as follows. Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible. Corollary 4.19 tells when the manifold $M$ with two specific balls removed is not contractible.
@@ 2. Preliminaries
This section is devoted to introducing some definitions and properties of finite posets and simplicial complexes that are needed in the sequel. We refer the reader to [1, 10, 14, 16, 17] for further details or proofs.
@@ 2.1. Finite posets
A topological space is said to be Alexandroff if every intersection of open sets is still an open set. If $X$ is an Alexandroff space, for each $x\in X,$ we denote by $U_{x}$ the intersection of all the open sets of $X$ containing $x$. This allows us to define a reflexive and transitive relation on $X$ by declaring $x\leq y$ if $U_{x}\subseteq U_{y}.$ The relation is antisymmetric if and only if $X$ is $T_{0}$. On the other hand, if $X$ is a set endowed with a reflexive, transitive relation, then the sets $U_{x}=\{y\in X|y\leq x\}$, as a basis, determines an Alexandroff topology. The Alexandroff topology is $T_{0}$ if and only if the relation is antisymmetric, that is, a partial order. A set with a partial order is called a poset. It has been shown that Alexandroff $T_{0}$ spaces and posets are in bijective correspondence. A finite $T_{0}$ space is clearly an Alexandroff $T_{0}$ space, thus a finite $T_{0}$ space is equivalent to a finite poset.
@@
Given a poset $X$ and an element $x\in X$, we denote $U_{x}=\{y\in X~|~y\leq x\},~\widehat{U}_{x}=\{y\in X~|~y<x\},$ $F_{x}=\{y\in X~|~x\leq y\},$ and $\widehat{F}_{x}=\{y\in X~|~x<y\}.$ Let $C_{x}=U_{x}\cup F_{x}$ and $\widehat{C}_{x}=\widehat{U}_{x}\cup\widehat{F}_{x}$. A totally ordered subset $C$ of $X$ is called a chain. An $n$-chain of $X$ is a chain of cardinality $n+1$.
@@
A finite poset $X$ is homogeneous of dimension $n$ if all maximal chains in $X$ have cardinality $n+1$. In this case, the degree of $x$, denoted ${\rm deg}(x)$, is the dimension of $U_{x}$.
If $X$ is a poset, we define its order complex $\mathcal{K}(X)$ as the ordered simplicial complex whose simplices are the nonempty chains of $X$. Conversely, given a simplicial complex $K$, we define the face poset $\chi(K)$ of $K$ as the poset of simplices of $K$ ordered by inclusion.
@@
Henceforth, all maps between posets that we deal with are assumed to be order-preserving.
@@ 2.2. Simplicial complexes
Let $K$ be a simplicial complex. A simplex $A\subseteq K$ is called a maximal simplex if $A$ is not a proper face of any simplex of $K$. The face $B$ of $A$ is called a free face of $A$ in $K$ if $B$ is the proper face of no other simplex of $K$.
$K$ is a homogeneous $k$-complex if every maximal simplex of $K$ is $k$-dimensional.
@@
[…]
It is true that $K\searrow L$ implies $|K|\searrow|L|$, but the converse is unknown.
3. Discrete homotopy theory for finite posets
@@
Let $\mathbb{Z}$ denote the digraph whose vertices are integers and whose edges are the ordered pairs $(2i,2i+1)$ and (2j,2j1)(2j,2j-1). Given an integer $p\geq 0$, write $I_{p}$ for the subgraph of $\mathbb{Z}$ induced by $\{0,\cdots,p\}$.
@@
Let $\mathbb{Z}^{n}$ be the $n$-fold Cartesian product of $\mathbb{Z}$ whose vertices are $n$-tuples of integers. For any interval $J\subseteq\mathbb{R}$, $\mathbb{Z}^{n}_{J}$ denotes the subgraph consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}$ with $x_{i}\in J$. Let $I^{n}_{p}$ be the $n$-fold Cartesian product of $I_{p}$. Denote by $\partial I^{n}_{p}$ the subgraph of $I^{n}_{p}$ consisting of all vertices $x=(x_{1},\cdots,x_{n})\in I^{n}_{p}$ with $x_{i}=p$ or 0 for some $i$.
@@
if there exists an integer $r\geq 0$ such that $f$ is a map $f:(\mathbb{Z}^{n},\mathbb{Z}_{[r,+\infty]}^{n})\rightarrow(X,x_{0})$. The minimum of such $r$ is called the radius of $f$.
@@
For a finite poset $X$ with basepoint $x_{0}\in X$, define $\pi_{n}^{D}(X,x_{0})$ to be the set of basepoint-preserving-homotopy classes of maps $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$. We call $\pi_{n}^{D}(X,x_{0})$ the $n$-th discrete homotopy group of $X$.
@@
Let $L^{n}$ be the $n$-dimensional cube, the product of $n$ copies of the interval [1,1][-1,1]. For each $i$-th copy of $L^{n}$, consider a partition
(3.1) $1=t_{-{j_{i}}^{\prime}}<\cdots<t_{-1}<0=t_{0}<t_{1}<\cdots<t_{j_{i}}=1,$
where $j_{i}$$^{\prime}$ and $j_{i}$ are the numbers of partition points between [1,0)[-1,0) and $(0,1]$, respectively. Then we define a partition $T$ of $L_{n}$ by
(3.2) $T=\bigcup_{\begin{subarray}{c}t_{-{j_{i}}^{\prime}}\leq k\leq t_{j_{i}}\\ 1\leq i\leq n\end{subarray}}t_{k}.$
With this partition $T$, $L^{n}$ can be identified with the geometric realization of a subgraph $\gamma_{T}$ of $\mathbb{Z}^{n}$,
$\gamma_{T}=\prod_{i=1}^{n}\mathbb{Z}_{[-{j_{i}}^{\prime},j_{i}]}.$
For each vertex $(x_{1},x_{2},\cdots,x_{n})$ of $\gamma_{T}$, taking it to the point $(t_{x_{1}},t_{x_{2}},\cdots,t_{x_{n}})$ of $L^{n}$ induces a canonical homeomorphism
$\lambda_{T}:|\mathcal{K}(\gamma_{T})|$ $\longrightarrow L^{n}.$
When the partition $T$ takes the form
$1=\frac{-r}{r}<\frac{-r+1}{r}<\cdots<\frac{-1}{r}<0=t_{0}<\frac{1}{r}\cdots<\frac{r-1}{r}<\frac{r}{r}=1$
for each $i$, where $r>0$ is an integer, we will write $T$, $\gamma_{T}$ and $\lambda_{T}$ as $T_{r}$, $\gamma_{r}$ and $\lambda_{r}$, respectively.
@@
It is straightforward to check that $\theta$ is well-defined and is a homomorphism. The following is one of the main theorems of this paper.
@@
There is a certain amount of technical machinery to be set up in order to prove this result. Given a partition $T$ of $L^{n}$ defined as (3.2), write $L^{n}$ as $\lambda_{T}|\mathcal{K}(\gamma_{T})|$. To each continuous map $f:\lambda_{T}|\mathcal{K}(\gamma_{T})|\rightarrow|\mathcal{K}(X)|$ which takes every simplex of $\lambda_{T}|\mathcal{K}(\gamma_{T})|$ into some simplex of $|\mathcal{K}(X)|$, we associate a quotient space
$J^{f}_{T}(|\mathcal{K}(X)|)$
of $|\mathcal{K}(X)|$ by the following procedure. Consider two vertices
$b=\lambda_{T}(a),~b^{\prime}=\lambda_{T}(a^{\prime})$
in $\lambda_{T}|\mathcal{K}(\gamma_{T})|$ with $a<a^{\prime}$. Then there is a simplex in $|\mathcal{K}(X)|$ containing $f(b)$ and $f(b^{\prime})$, denoted $\sigma$. Let $l_{f(b)f(b^{\prime})}$ be a simple curve in $\sigma$ connecting $f(b)$ and $f(b^{\prime})$, which is disjoint from the images of all the other vertices of $\lambda_{T}|\mathcal{K}(\gamma_{T})|$ under $f$. Furthermore, $l_{f(b)f(b^{\prime})}$ can be chosen to satisfy that for every point $x$ in $l_{f(b)f(b^{\prime})}-\{f(b),f(b^{\prime})\}$,
(3.3) $\text{support}(x)=\begin{cases}\text{support}(f(b)),~\text{if}~\mu_{X}(f(b))\leq\mu_{X}(f(b^{\prime}))\\ \text{support}(f(b^{\prime})),~\text{if}~\mu_{X}(f(b^{\prime}))\leq\mu_{X}(f(b)).\end{cases}$
Let us form the quotient space $J^{f}_{T}(|\mathcal{K}(X)|)$ under the identifications
(3.4) $x\sim\begin{cases}f(b),~\text{if}~\text{support}(x)=\text{support}(f(b)),\\ f(b^{\prime}),~\text{if}~\text{support}(x)=\text{support}(f(b^{\prime})),\end{cases}$
where $x\in l_{f(b)f(b^{\prime})}-\{f(b),f(b^{\prime})\}$, and $b,b^{\prime}$ range over all the pairs which are the images under $\lambda_{T}$ of the comparable pairs in $\gamma_{T}$. It is easy to see that by construction $\mu_{X}$ is well-defined on the quotient space $J^{f}_{T}(|\mathcal{K}(X)|)$. We write $\widetilde{\mu}_{X}$ for this induced map.
Before proceeding further we need some technical lemmas about the above constructions.
Lemma 3.4.
$\widetilde{\mu}_{X}:J^{f}_{T}(|\mathcal{K}(X)|)\rightarrow X$ is continuous.
@@
$\mu_{X}^{-1}(U^{X}_{f(b^{\prime})})=|\mathcal{K}(X)|-|Z|.$
Note that (3.3) and (3.4) imply
$l_{f(b)f(b^{\prime})}\cap|Z|=\emptyset.$
[…]
Lemma 3.5.
The map
$J_{T}(f):\gamma_{T}\rightarrow J^{f}_{T}(|\mathcal{K}(X)|)$
defined by $J_{T}(f)(a)=[f(\lambda_{T}(a))]$ is continuous.
Proof.
We may assume that $n=1$ here, since the topology on $\gamma_{T}$ is given by product topology. Given a point $f(\lambda_{T}(a))]\in J^{f}_{T}(|\mathcal{K}(X)|)$, let $B_{[f(\lambda_{T}(a))]}\subseteq J^{f}_{T}(|\mathcal{K}(X)|)$ be a ball centered at $f(\lambda_{T}(a))]$ with radius sufficiently small such that
$\pi^{-1}(B_{[f(\lambda_{T}(a))]})\cap f(\gamma_{T})=\{f(\lambda_{T}(a))\}.$
We will prove that $J_{T}(f)^{-1}(B_{[f(\lambda_{T}(a))]})$ is open which will shows that $J_{T}(f)$ is continuous.
Let $a$ is an odd integer. We can assume without loss of generality that $U^{\gamma_{T}}_{a}=\{a-1,a,a+1\}$. If $\mu_{X}(f(\lambda_{T}(a-1)))\leq\mu_{X}(f(\lambda_{T}(a)))$, then $l_{f(\lambda_{T}(a-1))f(\lambda_{T}(a))}-\{f(\lambda_{T}(a))\}$ is identified with $f(\lambda_{T}(a-1))$. The fact
$\pi^{-1}(B_{[f(\lambda_{T}(a))]})\cap\{l_{f(\lambda_{T}(a-1))f(\lambda_{T}(a))}-\{f(\lambda_{T}(a))\}\}\not$$\emptyset$
says that
$a-1\in J_{T}(f)^{-1}(B_{[f(\lambda_{T}(a))]}).$
On the other hand, if $\mu_{X}(f(\lambda_{T}(a)))\leq\mu_{X}(f(\lambda_{T}(a-1)))$, then $l_{f(\lambda_{T}(a-1))f(\lambda_{T}(a))}-\{f(\lambda_{T}(a-1))\}$ is identified with $f(\lambda_{T}(a))$. For any point $b\in l_{f(\lambda_{T}(a-1))f(\lambda_{T}(a))}-\{f(\lambda_{T}(a-1))\}$, we have
$B_{[f(\lambda_{T}(a))]}=B_{[b]}.$
Note that $\pi^{-1}(B_{[b]})$ is an open ball containing $b$ in $|\mathcal{K}(X)|$. When $b$ becomes arbitrarily close to $f(\lambda_{T}(a-1))$, $f(\lambda_{T}(a-1))\in\pi^{-1}(B_{[b]})$. Then it follows that
$a-1\in J_{T}(f)^{-1}(B_{[b]})=J_{T}(f)^{-1}(B_{[f(\lambda_{T}(a))]}).$
The same argument proves that
$a+1\in J_{T}(f)^{-1}(B_{[f(\lambda_{T}(a))]}).$
Thus by the choice of $B_{[f(\lambda_{T}(a))]}$ we conclude that
$U^{\gamma_{T}}_{a}=J_{T}(f)^{-1}(B_{[f(\lambda_{T}(a))]})$
which is the open set in $\gamma_{T}$.
The case $a$ is an even integer can be handled by similar arguments. Hence $J_{T}(f)$ is continuous, as desired. ∎
Remark 3.6.
Composing the maps in the previous two lemmas yields an order-preserving map
$\widetilde{\mu}_{X}\circ J_{T}(f):\gamma_{T}\rightarrow X.$
An especially important fact is that when the domain of $f$ is pulled back to $\gamma_{T}$ by the homeomorphism $\lambda_{T}$, $\mu_{X}\circ f$ and $\widetilde{\mu}_{X}\circ J_{T}(f)$ take the same values on every element of $\gamma_{T}$, that is,
$\mu_{X}\circ f\circ\lambda_{T}|\gamma_{T}=\widetilde{\mu}_{X}\circ J_{T}(f).$
This allows $\widetilde{\mu}_{X}\circ J_{T}(f)$ to serve as a useful tool for the reconstruction and approximation of topological properties of $f$.
Hence, when $f\in\pi_{n}(|\mathcal{K}(X)|,x_{0})$, one might hope that $\widetilde{\mu}_{X}\circ J_{T}(f)$ could provide all necessary details to form a preimage of $f$ in $\pi_{n}^{D}(X,x_{0})$. This is indeed the case as we shall see.
To get a better understanding of the proof of this theorem, let us sketch the proof strategy. The main idea of proving this theorem is to pull back information about $X$ to $|\mathcal{K}(X)|$ via the powerful tool, the K-McCord map $\mu_{X}$. The injectivity part will not be difficult once we finish the proof of the surjectivity part. The harder part of the proof is to show $\theta$ is surjective. In order to achieve this, when given a map $f:L^{n}\rightarrow|\mathcal{K}(X)|$, we first reconstruct it in terms of a simpler one, a simplicial map $f^{\prime}$. As stated in Remark 3.6, we define a preimage $g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ in terms of $f^{\prime}$, namely,
$\theta([g])=[|\mathcal{K}(g)|\circ\lambda_{r_{g}}^{-1}]=[f].$
The fact that $\mu_{X}$ is a weak homotopy equivalence allows us to turn to prove
$\mu_{X}\circ\mathcal{K}(g)\circ\lambda^{-1}_{r_{g}}]=[{\mu_{X}}\circ|f].$
To prove this, a finite sequence of maps will be interpolated between them, each map in the sequence being comparable with its neighbors. We now start the proof of the theorem, filling in the technical details.
Proof of Theorem 3.3.
We first show that $\theta$ is surjective. Let $f\in\pi_{n}(|\mathcal{K}(X)|,x_{0})$. Using the simplicial approximation theorem [15, Theorem 16.1], we can represent $f$ by a simplicial map $f^{\prime}$ from $K$ to $\mathcal{K}(X)$, where $K$ is some iterated barycentric subdivision of $L^{n}$. The key observation now is that for an arbitrary integer $r>0$, under the homeomorphism $\lambda_{r}$, $\mathcal{K}(\gamma_{r})$ becomes a triangulation of $L^{n}$. When $r$ is sufficiently large, $\mathcal{K}(\gamma_{r})$ can be regarded as a refinement of $K$.
Given a simplex $\sigma$ of $\mathcal{K}(X)$, let $|\sigma|_{\epsilon}$ be the open set in $|\mathcal{K}(X)|$ with the following properties:
@@
    $|\sigma|\subseteq|\sigma|_{\epsilon}$.